QUESTION 18

  1. The labor force is the

    number of employed people plus the number of unemployed people.

    total population divided by the number of employed people.

    number of employed people in the working-age population.

    number of employed people minus the number of unemployed people.

2 points   

QUESTION 19

  1. Category

    Number of People

    (Millions)

    Total Population 300
    Working-Age Population 220
    Labor Force 145
    Employed Workers 140

    Calculate the unemployment rate. (Round to two decimal places)

    Calculate the Labor Force Participation Rate.  (Round to two decimal places)

4 points   

QUESTION 20

  1. The unemployment rate generally ________ during recessions and generally ________ during expansions.

    decreases; decreases

    increases; does not change

    increases; decreases

    decreases; increases
Answer #1

18. a. number of employed people plus the number of unemployed people.

The labor force is the labor pool in employment. Labor force= employed + unemployed.

19. Unemployed = labor force - employed = 145- 140 = 5.

Unemployment rate = unemployed/ labor force x 100 = 5/145 x 100 = 3.44%

Labor Force Participation Rate = labor force/ working age population x 100 = 145 /220 x 100 = 65.90%

20. c. increases; decreases

During recession the business activity reaches its minimum and thus a lot of workers are laid off. The unempoyment rate increases during recession and the opposite happens during expansion when businesses are booming.